Ex: a) For the three charges shown below, calculate the force on charge 2 (magnitude and direction) due to the other two charges <-0.30 m ----><--0.20 m---> ㄑㄧ Q1 Q2 5.0 uC Q3 -3.0 HC -3.0 uC b) What is the magnitude of the electric field at q2?
Answer #1

a)

F = k q2 q3 / r1^2 - k q1 q2 / r2^2

F = 9*10^9* 10^-12* 3*5 ( 1/ 0.2^ - 1/ 0.3^2)

F = 336 N

B)

Electric field, E = 336/(5 x 10^-6)= 6.72 x 10^7 N/C

Comment in case any doubt please rate my answer....